Please show me how to solve this problem. a+1/8=1/2
Answer by Here2Help(19)

a+1/8=1/2
a+1/8-1/8=1/2-1/8
a=3/8
check:
3/8+1/8=1/2
4/8=1/2
4/8 reduces to 1/2
Hope this helps